OpenStudy (kaylala):

Find the value of the following using the concepts of both addition and subtraction formulas and reference angle: sin 255

OpenStudy (kaylala):

@eliassaab

OpenStudy (kaylala):

@ganeshie8

OpenStudy (kaylala):

@Callisto

OpenStudy (anonymous):

\[ \sin (225)=-\sin (360-255)=-\sin (105)=-\sin (45+60)=\\-\sin (45)) \cos (60)-\sin (60) \cos (45) \]

OpenStudy (anonymous):

\[ -\sin (45)) \cos (60)-\sin (60) \cos (45)=-\frac{1+\sqrt{3}}{2 \sqrt{2}} \]

OpenStudy (kaylala):

why is it -sin(360-255) why the - sign there?

OpenStudy (anonymous):

\[ -\sin(360-255)=- \sin (-255) =-(-\sin( 255)) = \sin(255) \]

OpenStudy (anonymous):

Did you understand it?

OpenStudy (kaylala):

no sorry where did you get that?

OpenStudy (anonymous):

If you add 360 to an angle you do not change the sine of that angle sin(360+y) = sin(y) sin( 360 +(-x)) =sin(-x) = -sin(x) If you do not know these elementary facts, please review them in your book or notes.

OpenStudy (anonymous):

Have you heard about periodic functions?

OpenStudy (kaylala):

have no heard of periodic functions ever @eliassaab